OpenStudy (anonymous):

A wooden boxcar is moving along a straight railroad track at speed V1. A sniper fires a bullet (initial speed V2) at it from a high-powered rifle. The bullet passes through both lengthwise walls of the car, its entrance and exit holes being exactly opposite each other as viewed from within the car. From what direction, relative to the track, is the bullet fired? Assume that the bullet is not deflected upon entering the car, but that its speed decreases by 20%. Take Vi = 85 kmlh and V2 = 650 mls. (Why don't you need to know the width of the boxcar?)

OpenStudy (anonymous):

I didn’t get the main idea of the problem. And why don’t we need the width of the boxcar?

OpenStudy (anonymous):

the hit only occur in normal direction,conserve moviment quantity(Q=m.v)

OpenStudy (anonymous):

it's not a momentum problem.

OpenStudy (anonymous):

Yeah, Algebraic is right. its a kinematic problem.

OpenStudy (anonymous):

Oh I got it. your sketch is almost right. I think this happens because the car is moving while the bullet is passing the width of the car:  so: \[V_2tsin \theta=V_1t \rightarrow \sin \theta=\frac{ V_1 }{ V_2}\] after substituting we get θ=2.6° thanks for help!